40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1081     Entry time: Fri Oct 24 10:06:16 2008
Author: Yoichi 
Type: Configuration 
Category: IOO 
Subject: MC gain and FSS gain changed 
Following the measurements of MC and FSS loop gains, I modified mcup script to set the MC VCO gain to 2dB (it was -4dB before).
I also changed the normal value of the FSS common gain to 7dB. The open loop transfer functions posted in the previous two entries
were measured with those settings.
ELOG V3.1.3-